SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Aug 23, 2005)

Attendees: Kris Chaisanguanthum, Fang Fang, Selina Li, Brian Lindquist, Dan Walker.

Operations

  • SVT saw a slight shift in timing of about 16ns for runs 57156-57210 after PEP fixed the timing fiducial problme on last Monday (Aug.15,2005). A similar shift was also seen in DIRC. See HN posted. This timing shift stopped after Chris O'Grady reset the master crate on Aug. 16, 2005. Selina showed a plot of t0 for the 24 Trigger lines, comparing two runs before and after timing shift disappeared.
  • We got paged due to a red alarm for damaged events with high percentage last Thursday, but it turned out to be SVT or ODF damaged events.

EMT

  • Dan has been successfully tested createFileTable with the EMT prescale configuration from his own database. He is working on getting this tested in the test stand. Andy will check Dan's header file first.

DCT/DCZ

  • Ming handed Selina a folder with some TSF related stuff (plots, flash cards, etc.) last Friday before he left SLAC. Ming will be presenting a poster on BaBar drift chamber trigger at CERN.
  • Now that the deadtime due to GLT is compariable to that of EMT, Jamie brought up the issue of prescaling ZPD data. He mentioned that the ZPD prescaling is a simple configuration parameter in the database. We will discuss this issue more for next week's Trigger operation meeting when more people can attend.

GLT

  • Valerie and Selina put in the new GLT FC firmware at IR2 on Aug. 16. All runs from run 57211 and onward have the new firmware in. There are still a small fraction of events with GLT word corruptions and EMT/DCT->GLT transmission errors. See HN posted.

L3

  • Last Thursday, L3T damage alarm showed 100% damaged events in COSMICS_NODCH runs. Rainer updated the crate mask for this run type. See HN posted.
  • Selina has the L3 pager now since no one is taking it. Arik is helping with some L3 related tasks.

Data Quality and Monitoring

  • There are updated instruction for the GLT page in JAS.
  • Brian showed some fastmon time history plots for TSF segments (along with other event time plots produced by Su Dong) for a run (45393) from Feb. 2004 where DCH tripped. One question we have is why the end run time in the xtc file is not the same as that indicated in the elog?

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.